An investment broker has received $250,000 to invest in a 12-month commitment. The money can be placed in Treasury notes (with a return of 8% and a risk score of 2) or in municipal bonds (with a return of 9% and a risk score of 3). The broker’s client wants diversification to the extent that between 50% and 70% of the total investment must be placed in Treasury notes. Also, because of fear of default, the client requests that the average risk score of the total investment should be no more than 2.42. How much should the broker invest in each security so as to maximize return on investment?
Answer #1

Investment made in treasury notes = 175,000

Investment made in municipal bonds = 75,000

Maximum return = 20750

Dole Poge Let hunted Let het investment made in treasury notes = x Let investment made in municipal bonds We must maximize thThere fore Max. Z=0.08xt +0.094 Subject to : xty £250,000 x 20.7Cty) 2x + 3y - 2.42 Caty) sys Find points of equations : Х 23Date Dage 26 (itty - 777250,000 - dok D in @ = 0,7(250,000 -y+y) y = 75,000 in Put value of a 250000-y from Put value of y atDole Page 105.000 + x = 3250000 2- 145000 Point of intersection = (145000, 105000) Hook book csak ery Hook (25олоо x=0.5% + 1Data Page Value of 2 Conner Points 0 (0,0) IA (145 000, 105000) B C175000, 75000) 6:08 xo 40.0480 -0. 0.08x145000 +0.09x 10 S